# Activation Addition: Steering Language Models Without Optimization

### v4

@@ Appendix A Broader Impacts
As the examples of anger- and conspiracy-steering show (Appendix Table 6), ActAdd  can easily be misused. Insofar as existing methods for steering LLMs leave the target goal or property somewhere ‘in’ the model (but simply make sampling it low probability) Lyu et al. [2024], activation engineering may circumvent superficial alignment methods.
@@ Appendix B Is ActAdd just a subtle kind of prompt engineering?
Generation scoring score ActAdd generations over different injection layers GPT-2-XL wedding N/A Fig7,8 Link
@@
Topic steering examples of topic control GPT-2-XL Various N/A Fig7,2 Link
Inference cost to inference GPT-2 This is a N/A Fig4 Link
overhead speed over increasing model size OPT test prompt
@@
Other run some of the GPT-J-6B Various N/A Fig12,13,14 Link
replications above on other models LLaMA-13B Tab13
Table 6: Note: some completions contain unpleasant content including slurs. Example GPT-2-XL outputs before and after applying ActAdd. A runnable notebook showing the best-of-3 samples used can be found here: tinyurl.com/ActAddExamples.
@@
Figure 6: Performance of ActAdd  on a target topic as the topic becomes more relevant. The perplexity ratio (lower better) compares the relative predictive performance of ActAdd and an unmodified model; we see that adding a wedding - ‘   ’ steering vector improves performance on wedding-related text while preserving performance on unrelated text.
@@
Figure 7: Topic steering effect (mean related words in completions) over injection layer. In blue is the average related-word count among 200 ActAdd  completions; the dotted line is the count among unmodified completions. Model: GPT-2-XL.
Figure 8: Topic steering effect (probability of seeing some related words in completions) over injection layer. In blue is the average probability among 200 ActAdd  completions; dotted line is the probability in unmodified completions. Model: GPT-2-XL.
@@ Appendix C Implementation details
To understand the event “Paul went to a vegetarian restaurant.”, it is important to know that vegetarian restaurants do not serve ’ meat
@@
Table 11: Tokens with the greatest absolute change in log probability under ActAdd(weddings). (See Figure 9 for the distribution these are drawn from.) The probabilities most increased on average are primarily wedding-related, with the exception of ‘OG’ and ‘08’. (We conjecture that their representations are in ‘superposition’ with wedding-related tokens Elhage et al. [2022]). The bottom tokens share no obvious theme and show a significantly lower absolute change in probability: the mean log-prob diff for token ‘ bride’ represents a probability increase of 500%, whereas for ‘Image’ it’s -30%.
@@ Appendix D Advantages of ActAdd over fine-tuning
Future work could compare ActAdd on knowledge editing benchmarks (Wu et al. [2023], Zhang et al. [2024]. (This comparison could be unfair to methods like ROME Meng et al. [2023]), since our method is not editing weights, but it would give standardized evidence about steering.)
Appendix E Reproducibility
The following represents an exhaustive list of models used, sampling strategies used, and searches run.
Models
We use GPT-2-XL (Radford et al. [2019]) (Modified MIT License) because GPT-2-small and GPT-2-medium were not capable enough to demonstrate the method: it was hard to tell whether the activation additions were failing or the model was failing. GPT-2-XL was the third language model we tried.
After observing success with GPT-2-XL, to replicate our results, we subsequently repeated the same experiments with Llama-1-13B Touvron et al. [2023] (open noncommercial nonmilitary license) and GPT-J-6B Wang and Komatsuzaki [2021] (apache-2.0 license). Our toxicity and sentiment experiments use OPT (Zhang et al. [2022b]) (open noncommercial nonmilitary license) and LLaMA-3-8B Meta [2024] (freely licensed for $<700$ million monthly active users). See Appendix F for details.
We use all-MiniLM-L6-v2 (Reimers and Gurevych [2019]) to compute sentence embeddings to calculate relevance using cosine similarity. For the success score, we use the SiEBERT (Hartmann et al. [2023b]) sentiment classifier.
We perform sentiment classificaton with the SiEBERT classifier Hartmann et al. [2023a].
APIs
For scoring toxicity, we use the Perspective API: perspectiveapi.com.
For scoring fluency, we use OpenAI davinci-002. The PREADD baseline instead used the discontinued davinci-[001] model.
Datasets
The Stanford IMDb Large Movie Review Dataset ([Maas et al., 2011]). No license: ‘we release this dataset to the public’ by the authors.
RealToxicityPrompts ([Gehman et al., 2020]). Apache 2.0.
LAMA ConceptNet ([Petroni et al., 2019]). CC-BY-NC 4.0.
OpenWebText ([Peterson et al., 2018]). GPL-3.0.
Seed
Sampling hyperparameters
We precommitted to fixed sampling hyperparameters, selected before experiments began. We held them fixed throughout our data collection. Those sampling hyperparameters were temperature$=1.0$, freq_penalty$=1.0$, and top_p=$0.3$. Since this top_p value seemed a bit unusual to us in retrospect, we invited [anonymized] to reproduce this process with an unmodified GPT-2-XL and report the best sampling hyperparameters they found. This second experiment was blinded, as they did not know the values we used. They found that temperature$=0.6$ and top_p$=0.5$ produced better GPT-2-XL capabilities. We reran all our qualitative results at this setting, and they all reproduced (subjectively, more impressively).
In replicating the unsteered OPT sentiment baseline, we find that the NegToPos direction is consistently higher success than PosToNeg. This holds across different combinations of model hyperparameters, including those in Pei et al. 2023. However, PREADD [Pei et al., 2023] reports similar success results for both (i.e. a much lower NegToPos success). We have used our calculated values for the OPT results.
Reporting the best of $K$ completions
We generated $K=3$ completions for each demonstration, for both normal and steered forward-passes. Appendix Table 6, shows the subjectively most compelling completion pair out of the first three seed-00 completion-pairs. You can see all top-3 completions for the entries in this notebook: tinyurl.com/actadd3.
We share activation additions which work well. We iterated over contrast pairs to get these to work, although several striking results were generated within [first author’s] first hour of using the technique.
Out of the 12 activation additions we thought demonstrated a distinct ability of the method, we decided not to include 1 because its first three seed-00 completions were unusually unimpressive. We include the remaining 11 in Table 6 and the above notebook.
ActAdd  hyperparameters ($l,c$)
Hardware:
  • GPU: Nvidia RTX A5000
  • CPU: Intel Core i9-10900X CPU @ 3.70GHz
Memory:
  • 24GB GPU RAM
  • 32GB system RAM
Names and versions of relevant libraries and frameworks:
  • Operating system: Ubuntu 22.04.1 LTS
  • numpy: 1.24.3
  • pandas: 2.0.1
  • torch: 1.13.1
  • transformer-lens: 1.4.0
@@ Appendix F Replicability
F.1 GPT-J-6B
Figures 12, 13, and 14 show the results from repeating the main experiments on GPT-J-6B Wang and Komatsuzaki [2021]. We see the same dynamics from the wedding vector running example: a targeted effect on only wedding-related tokens (using both KL-div and token probability); and similar effects when injected at different layers of GPT-J and with different magnitudes $c$ applied.
F.2 Llama-1-13B
Table 13 sees ActAdd  displaying the same qualitative steering effect when applied to Llama-1-13B Touvron et al. [2023] (though with a notable failure to replicate on Example 6, Paris $\rightarrow$ Rome, the anger vector, and the harm vector). [^2] See this runnable notebook for the full Llama-1-13B run using $k=3$ sampling; the output is the result from the first and only execution: tinyurl.com/actadd3-llama.
F.3 OPT-6.7B
We use the OPT model Zhang et al. [2022b] in our toxicity (Table 3) and sentiment (Table 4) experiments. ActAdd-OPT using the love--hate vector produces a statistically significant 17% drop in toxicity over an unsteered OPT, at a small (partially unavoidable owing to the nature of the detoxification task) cost to fluency and relevance. ActAdd-OPT using the love--hate vector produces a 21% absolute increase in positive classification over an unsteered OPT, at a larger (partially unavoidable owing to the nature of the sentiment shift task) cost to fluency and relevance.
F.4 Llama-3-8B
We also use Llama-3-8B Meta [2024] in our toxicity (Table 3) and sentiment (Table 4) experiments. ActAdd-LLaMA-3 using the love--hate vector produces a statistically significant 5% drop in toxicity over an unsteered Llama-3-8B, at a very small (partially unavoidable owing to the nature of the detoxification task) cost to fluency and relevance. ActAdd-LLaMA-3 using the love--hate vector produces a 25% absolute increase in negative-to-positive classification over an unsteered Llama-3-8B, at a larger (partially unavoidable owing to the nature of the sentiment shift task) cost to fluency and relevance.
Appendix G Investigating the norm of steering vectors
@@
Appendix H Investigating random ActAdd vectors
@@
Appendix I Partial ActAdd
@@
Table 13: Example outputs before and after applying ActAdd on Llama-13B. Notice the failed replication on the Eiffel example. A notebook showing the best-of-3 samples used can be found here: tinyurl.com/actadd3-llama.
@@
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: See Section 4
  5. 2.
    Limitations
  6. Question: Does the paper discuss the limitations of the work performed by the authors?
  7. Answer: [Yes]
  8. Justification: Section 5 and Appendix C
  9. 3.
    Theory Assumptions and Proofs
  10. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  11. Answer: [N/A]
  12. Justification: N/A
  13. 4.
    Experimental Result Reproducibility
  14. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  15. Answer: [Yes]
  16. Justification: As well as Section 3, Appendices E,  F and C, all code is permanently archived at Zenodo. Most experiments can be run in the Google Colab environment, and links to runnable notebooks are provided for these experiments.
  17. 5.
    Open access to data and code
  18. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  19. Answer: [Yes]
  20. Justification: All datasets are open from other work. As well as Section 3, Appendices E,  F and C, all code is permanently archived at Zenodo. Most experiments can be run in the Google Colab environment, and links to runnable notebooks are provided for these experiments.
  21. 6.
    Experimental Setting/Details
  22. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  23. Answer: [Yes]
  24. Justification: Section 3, Appendices E,  F and C.
  25. 7.
    Experiment Statistical Significance
  26. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  27. Answer: [Yes]
  28. Justification: The toxicity and sentiment results all pass a two-sample $t$-test against the second-best method.
  29. 8.
    Experiments Compute Resources
  30. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  31. Answer: [Yes]
  32. Justification: Appendix E
  33. 9.
    Code Of Ethics
  34. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  35. Answer: [Yes]
  36. Justification: Fair Wages (N/A), Research involving human participants (N/A), Privacy (Yes), Consent (Yes), Deprecated datasets (Yes), Copyright and Fair Use (Yes), Representative evaluation practice (Yes), Safety (Yes), Security (Yes), Discrimination (Yes), Surveillance (Yes), Deception & Harassment (Yes), Environment (Yes), Human Rights (Yes), Bias and fairness (Yes), Data and model documentation (Yes), Data and model licenses (Yes), Secure and privacy-preserving data storage & distribution (N/A), Responsible release and publication strategy (Yes), Allowing access to research artifacts (Yes), Disclose essential elements for reproducibility (Yes), Ensure legal compliance (Yes).
  37. 10.
    Broader Impacts
  38. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  39. Answer: [Yes]
  40. Justification: See Appendix A.
  41. 11.
    Safeguards
  42. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  43. Answer: [N/A]
  44. Justification: No new data released, no new models released. See Appendix A for a discussion of the method’s risks.
  45. 12.
    Licenses for existing assets
  46. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  47. Answer: [Yes]
  48. Justification: See Appendix E.
  49. 13.
    New Assets
  50. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  51. Answer: [Yes]
  52. Justification: Main repo includes documentation.
  53. 14.
    Crowdsourcing and Research with Human Subjects
  54. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  55. Answer: [N/A]
  56. Justification: N/A
  57. 15.
    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
  58. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  59. Answer: [N/A]
  60. Justification: N/A

### v5

@@ Appendix A Broader Impacts
As the examples of anger- and conspiracy-steering show (Appendix Table 6), ActAdd  can easily be misused. Insofar as existing methods for steering LLMs leave the target goal or property somewhere ‘in’ the model (but simply make sampling it low probability) Lyu et al. 2024, activation engineering may circumvent superficial alignment methods.
@@ Appendix B Is ActAdd just a subtle kind of prompt engineering?
Generation scoring score ActAdd generations over different injection layers GPT-2-XL wedding N/A Fig7,3 Link
@@
Topic steering examples of topic control GPT-2-XL Various N/A Fig7,4 Link
@@
Table 6: Note: some completions contain unpleasant content including slurs. Example GPT-2-XL outputs before and after applying ActAdd.
@@
B.0.1 Experiment: Steering towards wedding topics
For this experiment, we use the following settings: $p^{*}=\mathrm{`I\,went\,up\,to\,my\,friend\,and\,said}\textit{'},p^{+}=\mathrm{`weddings}\textit{'},$ ’ ,
$p_{-}=\mathrm{`\,\textit{'}},c=1.0,\,\mathrm{seed}=0.$ Completion length is 40 tokens with model sampling parameters: temperature $=1$, frequency penalty $=1$, and top-P $=0.3$.
Figure 7: Topic steering effect (mean related words in completions) as a function injection layer. In blue is the average related-word count among 200 ActAdd  completions. The dotted line is the rate for the unsteered GPT-2-XL.
@@ Appendix C Implementation details
To understand the event “Paul went to a vegetarian restaurant.”, it is important to know that vegetarian restaurants do not serve meat
@@
Table 11: Tokens with the greatest absolute change in log probability under ActAdd(weddings). (See Figure 8 for the distribution these are drawn from.) The probabilities most increased on average are primarily wedding-related, with the exception of ‘OG’ and ‘08’. (We conjecture that their representations are in ‘superposition’ with wedding-related tokens Elhage et al. 2022). The bottom tokens share no obvious theme and show a significantly lower absolute change in probability: the mean log-prob diff for token ‘ bride’ represents a probability increase of 500%, whereas for ‘Image’ it’s -30%.
@@
C.1 Details of perplexity experiments
For each sentence in each document, we calculate the log-probabilities $\mathcal{L}(t_{k})$ for each token $t_{k}\in s_{j}$ under the unmodified $M_{\mathrm{baseline}}$ and modified $M_{\mathrm{\text{ActAdd}}}$ models.
[…]
@@ Appendix D Advantages of ActAdd over fine-tuning
Future work could compare ActAdd on knowledge editing benchmarks (Wu et al. 2023; Zhang et al. 2024. (This comparison could be unfair to methods like ROME Meng et al. 2023), since our method is not editing weights, but it would give standardized evidence about steering.)
@@ Appendix E Replicability
E.1 GPT-J-6B
Figures 11, 12, and 13 show the results from repeating the main experiments on GPT-J-6B Wang & Komatsuzaki 2021. We see the same dynamics from the wedding vector running example: a targeted effect on only wedding-related tokens (using both KL-div and token probability); and similar effects when injected at different layers of GPT-J and with different magnitudes $c$ applied.
E.2 Llama-1-13B
Table 13 sees ActAdd  displaying the same qualitative steering effect when applied to Llama-1-13B Touvron et al. 2023 (though with a notable failure to replicate on Example 6, Paris $\rightarrow$ Rome, the anger vector, and the harm vector).
E.3 OPT-6.7B
We use the OPT model Zhang et al. 2022b in our toxicity (Table 3) and sentiment (Table 4) experiments. ActAdd-OPT using the love--hate vector produces a statistically significant 17% drop in toxicity over an unsteered OPT, at a small (partially unavoidable owing to the nature of the detoxification task) cost to fluency and relevance. ActAdd-OPT using the love--hate vector produces a 21% absolute increase in positive classification over an unsteered OPT, at a larger (partially unavoidable owing to the nature of the sentiment shift task) cost to fluency and relevance.
E.4 Llama-3-8B
We also use Llama-3-8B Meta 2024 in our toxicity (Table 3) and sentiment (Table 4) experiments. ActAdd-LLaMA-3 using the love--hate vector produces a statistically significant 5% drop in toxicity over an unsteered Llama-3-8B, at a very small (partially unavoidable owing to the nature of the detoxification task) cost to fluency and relevance. ActAdd-LLaMA-3 using the love--hate vector produces a 25% absolute increase in negative-to-positive classification over an unsteered Llama-3-8B, at a larger (partially unavoidable owing to the nature of the sentiment shift task) cost to fluency and relevance.
Appendix F Investigating the norm of steering vectors
@@
Appendix G Investigating random ActAdd vectors
@@
Appendix H Partial ActAdd
@@
Table 13: Example outputs before and after applying ActAdd on Llama-13B. Notice the failed replication on the Eiffel example.